find the sum of 3 consecutive odd integers such that 4 times the sum of the 1st and second is 17 more than 7 times the 3rd: find the sum of 3 consecutive odd integers such that 4 times the sum of the 1st and second is 17 more than 7 times the 3rd
Answer by solver91311(1877)

First odd integer: x
Second consecutive odd integer: x+2
Third consecutive odd integer: x+4

The sum of the first and second: x + (x + 2)=2x +2
4 times the sum of the 1st and 2nd: 4(2x+2)=8x+8

7 times the third: 7(x+4)=7x+28
17 more than 7 times the 3rd: 7x+28+17=7x+45

So: 8x+8=7x+45. Solve this for x and verify that x is odd.

The sum of the three integers is: x + (x+2) + (x+4)=3x+6. Take the solution you found above, multiply by 3 and then add 6 to get your final answer.